# Characteristic Learning for Provable One Step Generation

### v3

@@ Appendix G Generalization Error Analysis for Least-Squares Regression
In this section, we provide a generalization error analysis for nonparametric regression, which is used in establishing oracle inequalities for velocity matching (Lemma D.1) and characteristic fitting (Lemma F.2).
@@
Lemma G.1.
Suppose that $|f^{*}(x)|\leq B$ for each $x\in{\mathcal{X}}$. Let ${\mathscr{F}}$ be a set of functions mapping from ${\mathcal{X}}$ to $[-B,B]$. Then it follows that for each $n\geq\operatorname{VCdim}({\mathscr{F}})$,
@@
This lemma provides a generalization error bound with fast rate via the technique of the offset Rademacher complexity, which was first proposed by Liang et al. [2015]. In recent years, this technique has been applied to the convergence rate analysis for deep nonparametric regression, such as Duan et al. [2023, Lemma 14] and Ding et al. [2024, Lemma 4.1].
Proof of Lemma G.1.
@@
=${\mathscr{H}}=\Big{\{}x\mapsto h(x)=(f(x)-f^{*}(x))^{2}:f\in{\mathscr{F}}\Big{\}}.$
@@
(G.1) ${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in{\mathscr{F}}}R(\widehat{f})-2\widehat{R}_{n}(\widehat{f})\Big{]}\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\xi}\Big{[}\sup_{h\in{\mathscr{H}}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h^{2}(X^{(i)})\Big{]}.$
We next estimate the expectation in the right-hand side of (G.1). Let $\delta\in(0,4B^{2})$ and ${\mathscr{H}}_{\delta}$ be a $L^{\infty}({\mathcal{D}})$ $\delta$-cover of ${\mathscr{H}}$ satisfying $|{\mathscr{H}}_{\delta}|=N(\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))$. Then for each $h\in{\mathscr{H}}$, there exists $h_{\delta}\in{\mathscr{H}}_{\delta}$ such that
@@
$-\frac{1}{n}\sum_{i=1}^{n}h^{2}(X^{(i)})+\frac{1}{n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\leq 8B^{2}\delta.$
@@
(G.2) $\leq{\mathbb{E}}_{\xi}\Big{[}\max_{h_{\delta}\in{\mathscr{H}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\Big{]}+5\delta.$
@@
$-h_{\delta}(X^{(i)})\leq\xi^{(i)}h_{\delta}(X^{(i)})\leq h_{\delta}(X^{(i)}),\quad 1\leq i\leq n.$
@@
(G.3) $\leq\exp\Big{(}-\frac{(\frac{nt}{3}+\frac{1}{12B^{2}}\sum_{i=1}^{n}h_{\delta}^{2}(t^{(i)},X_{t}^{(i)}))^{2}}{2\sum_{i=1}^{n}h_{\delta}^{2}(t^{(i)},X_{t}^{(i)})}\Big{)}\leq\exp\Big{(}-\frac{nt}{18B^{2}}\Big{)},$
@@
where the inequality is owing to (G.3). Letting $A=18B^{2}\log|{\mathscr{H}}_{\delta}|n^{-1}$ gives that
(G.4) ${\mathbb{E}}_{\xi}\Big{[}\max_{h_{\delta}\in{\mathscr{H}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\Big{]}\leq 18B^{2}\frac{\log|{\mathscr{H}}_{\delta}|+1}{n}.$
@@
(G.5) $\log N(\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))\leq\log N\Big{(}\frac{\delta}{4B},{\mathscr{F}},L^{\infty}({\mathcal{D}})\Big{)}\leq\operatorname{VCdim}({\mathscr{F}})\log\Big{(}\frac{4eB^{2}n}{\delta}\Big{)},$
where the first and last inequalities follows from Lemmas A.10 and A.11, respectively. Combining (G.1), (G.2), (G.4) and (G.5), we have
${\mathbb{E}}_{{\mathcal{D}}}\Big{[}\sup_{f\in{\mathscr{F}}}R(\widehat{f})-2\widehat{R}_{n}(\widehat{f})\Big{]}$ $\leq\inf_{\delta>0}\Big{\{}36B^{2}\frac{\operatorname{VCdim}({\mathscr{F}})}{n}\log\Big{(}\frac{4eB^{2}n}{\delta}\Big{)}+5\delta\Big{\}}.$
@@
Lemma G.2.
Suppose that $|f^{*}(x)|\leq B$ for each $x\in{\mathcal{X}}$. Let ${\mathscr{F}}$ be a set of functions mapping from ${\mathcal{X}}$ to $[-B,B]$. Let $\{\varepsilon^{(i)}\}_{i=1}^{n}$ be a set of independent $\sigma^{2}$-sub-Gaussian random variables. Then it follows that for each $n\geq\operatorname{VCdim}({\mathscr{F}})$,
@@
This proof uses a technique similar to the proof of Schmidt-Hieber [2020, Lemma 4].
Proof of Lemma G.2.
@@
(G.6) ${\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\varepsilon^{(i)}(\widehat{f}(X^{(i)})-\widehat{f}_{\delta}(X^{(i)}))\Big{]}\leq\delta{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\frac{1}{n}\sum_{i=1}^{n}|\varepsilon^{(i)}|\Big{]}\leq\delta\sigma,$
@@
(G.7) $\widehat{R}_{n}^{1/2}(\widehat{f}_{\delta})-\widehat{R}_{n}^{1/2}(\widehat{f})\leq\Big{(}\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{\delta}(X^{(i)})-\widehat{f}(X^{(i)}))^{2}\Big{)}^{1/2}\leq\delta.$
@@
(G.8) $\leq\frac{1}{4}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big{[}\widehat{R}_{n}(\widehat{f})\Big{]}+\frac{2}{n}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\big{[}\psi^{2}(\widehat{f}_{\delta})\big{]}+\frac{1}{4}\delta^{2}+\delta\sigma,$
@@
Here the first inequality follows from (G.6), the second inequality holds from (G.7), the third inequality is due to Cauchy-Schwarz inequality, and the last inequality is owing to the weighted AM-GM inequality $ab\leq a/4+b$ for each $a,b\in{\mathbb{R}}$.
@@
(G.9) ${\mathbb{E}}_{\varepsilon}\big{[}\psi^{2}(\widehat{f}_{\delta})\big{]}\leq{\mathbb{E}}_{\varepsilon}\Big{[}\max_{f_{\delta}\in{\mathscr{F}}_{\delta}}\psi^{2}(f_{\delta})\Big{]}\leq 4\sigma^{2}(\log|{\mathscr{F}}_{\delta}|+1).$
We now estimate the covering number $|{\mathscr{F}}_{\delta}|=N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))$. It follows from Lemma A.11 that for $n\geq\operatorname{VCdim}({\mathscr{F}})$,
(G.10) $\log N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))\leq\operatorname{VCdim}({\mathscr{F}})\log\Big{(}\frac{eBn}{\delta}\Big{)}.$
Combining (G.8), (G.9) and (G.10), and setting $\delta=B/n$ complete the proof. ∎
Appendix H Approximation by Deep Neural Networks with Lipschitz Constraint
The approximation error analysis for deep neural networks has been investigated by Yarotsky [2017, 2018], Yarotsky and Zhevnerchuk [2020], Shen et al. [2019], Shen [2020], Lu et al. [2021], Petersen and Voigtlaender [2018], Jiao et al. [2023a], Duan et al. [2022]. However, limited work has been done for deep neural networks with Lipschitz constraint [Huang et al., 2022, Chen et al., 2022, Jiao et al., 2023b, Ding et al., 2024].
This proof is based on the proof of Yarotsky [2017, Theorem 1]. The ReLU activation function is defined as $\operatorname{ReLU}(x)=\max\{0,x\}$, and the ReQU activation function is defined as the squared ReLU function $\operatorname{ReQU}(x)=(\max\{0,x\})^{2}$.
Lemma H.1.
@@ Appendix H Approximation by Deep Neural Networks with Lipschitz Constraint
Proof of Lemma H.2.
According to Li et al. [2019, Lemma 2.1], the following identities hold:
@@
Proof of Lemma H.3.
Define the augmented input vector $(x_{1},\ldots,x_{p},x_{p+1},\ldots,x_{n})$ where $n=2^{\lceil\log_{2}p\rceil}$ and $x_{i}=1$ for $p+1\leq i\leq n$. Observe that $\prod_{i=1}^{p}x_{i}=\prod_{i=1}^{n}x_{i}$. According to Lemma H.2, the mapping $(x_{1},\ldots,x_{n})\rightarrow(x_{1}x_{2},\ldots,x_{n-1}x_{n})\in{\mathbb{R}}^{n/2}$ can be implemented by a ReQU neural network with $1$ hidden layer and $6n$ non-zero parameters. By a same argument, we can construct a ReQU neural network with $1$ hidden layer and $3n$ non-zero parameters, which maps $(x_{1}x_{2},\ldots,x_{n-1}x_{n})$ to $(x_{1}x_{2}x_{3}x_{4},\ldots,x_{n-3}x_{n-2}x_{n-1}x_{n})\in{\mathbb{R}}^{n/4}$. By induction on $n$, the number of layers is given as $\lceil\log_{2}p\rceil$, and the total number of non-zero parameters is given by
@@
(H.1) $\psi(z)=\left\{\begin{aligned} &1,&&|z|<1,\\ &2-|z|,&&1\leq|z|\leq 2,\\ &0,&&2<|z|,\end{aligned}\right.$
@@
Noticing that $\widetilde{\psi}(z)=\min\{\operatorname{ReLU}(z+2),\operatorname{ReLU}(-z+2)\}$, by applying Lemma H.1, we find that $\widetilde{\psi}$ can be implemented by a ReLU neural network with $2$ hidden layers and $11$ non-zero parameters. Further, according to the equality $\psi(z)=\min\{1,\widetilde{\psi}(z)\}$, the univariate trapezoid function $\psi$ can be implemented by a ReLU neural network with $3$ hidden layers and $14$ non-zero parameters. This completes the proof. ∎
Lemma H.5 (Approximation error).
@@
Proof of Lemma H.5.
The proof is divided into three steps. In the first step, we approximate the target function based on a partition of unity and the degree-00 Taylor expansion. Then we implement this piece-wise linear function using deep neural network exactly in the second step. Finally, in the last step, we estimate the Lipschitz constant of the deep neural network.
@@
Consider a partition of unity formed by a grid of $\prod_{k=1}^{p}(N_{k}+1)$ functions $\phi_{m}$ on the domain $[0,1]^{p}$:
(H.2) $\sum_{m}\phi_{m}(x)\equiv 1,\quad x\in[0,1]^{p},$
@@
(H.3) $\phi_{m}(x)=\prod_{k=1}^{p}\psi\Big{(}3N_{k}\Big{(}x_{k}-\frac{m_{k}}{N_{k}}\Big{)}\Big{)}.$
Here $\psi$ is the univariate trapezoid function defined as (H.1). It is noticeable that for each $m$,
(H.4) $\sup_{z\in[0,1]}|\psi(z)|=1,\quad\sup_{x\in[0,1]^{p}}|\phi_{m}(x)|=1,$
@@
[…]
@@
(H.6) $u(x)=\sum_{m}\phi_{m}(x)u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}.$
@@
$\leq\sum_{m}\Big{|}u^{*}(x)-u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}\Big{|}\mathbbm{1}\Big{\{}m:\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\leq\frac{2}{3N_{k}},\leavevmode\nobreak\ k\in[p]\Big{\}}$
$\leq 2^{p}\max_{m}\Big{|}u^{*}(x)-u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}\Big{|}\mathbbm{1}\Big{\{}m:\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\leq\frac{2}{3N_{k}},\leavevmode\nobreak\ k\in[p]\Big{\}}$
$\leq 2^{p}\max_{m}\sum_{k=1}^{p}\mathop{\operatorname{ess}\,\sup}_{x\in[0,1]^{p}}|\partial_{k}u^{*}(x)|\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\mathbbm{1}\Big{\{}m:\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\leq\frac{2}{3N_{k}},\leavevmode\nobreak\ k\in[p]\Big{\}}$
@@
where the first inequality follows from the triangular inequality, the second inequality is due to (H.4) and (H.5), the third inequality used the observation that any $x\in[0,1]^{p}$ belongs to the support of at most $2^{d}$ functions $\phi_{m}$, the forth inequality used Taylor’s theorem of degree-00, and the last inequality holds from Hölder’s inequality. Consequently, we obtain the following inequality
@@
In this step, we implement the piececwise linear approximation (H.6) by a deep neural network. Using Lemmas H.3 and H.4, for each $m$, the function $\phi_{m}$ defined as (H.3) can be implemented by a deep neural network with $\lceil\log_{2}p\rceil+3$ layers and $616p+6(p+1)=22p+$ 6 non-zero parameters. Since $u$ defined in (H.6) is a linear combination of $\prod_{k=1}^{p}(N_{k}+1)$ functions $\phi_{m}$, it can be implemented by a deep neural network with $\lceil\log_{2}p\rceil+3$ layers and $(22p+6)\prod_{k=1}^{p}(N_{k}+1)$ non-zero parameters.
@@
According to Step 2, the piecewise linear approximation $u$ can be implemented by a deep neural network with no error. Therefore, it suffices to compute the Lipschitz constant of $u$ in (H.6). Taking derivative on both sides of (H.6) with respect to $x_{k}$ yields
@@
(H.7) $=\sum_{m}u^{*}\Big{(}\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big{)}A_{k}(m)\partial_{k}\psi\Big{(}3N_{k}\Big{(}x_{k}-\frac{m_{k}}{N_{k}}\Big{)}\Big{)}\mathbbm{1}\Big{\{}m:\Big{|}x_{k}-\frac{m_{k}}{N_{k}}\Big{|}\leq\frac{2}{3N_{k}}\Big{\}},$

### v5

@@ Appendix I Generalization Error Analysis for Least-Squares Regression
In this section, we provide a generalization error analysis for nonparametric regression, which is used in establishing oracle inequalities for velocity matching (Lemma E.1) and characteristic fitting (Lemma G.2).
@@
Lemma I.1.
Suppose that $|f^{*}(x)|\leq B$ for each $x\in{\mathcal{X}}$. Let $\mathscr{F}$} be a set of functions mapping from $\mathcal{X}$} to [B,B][-B,B]. Then it follows that for each $n\geq\operatorname{VCdim}({\mathscr{F}})$,
@@
This lemma provides a generalization error bound with fast rate via the technique of the offset Rademacher complexity, which was first proposed by Liang et al. [2015]. In recent years, this technique has been applied to the convergence rate analysis for deep nonparametric regression, such as [Duan et al., 2023, Lemma 14] and [Ding et al., 2024, Lemma 4.1].
Proof of Lemma I.1.
@@
${\mathscr{H}}=\Big\{x\mapsto h(x)=(f(x)-f^{*}(x))^{2}:f\in{\mathscr{F}}\Big\}.$
@@
(I.1) ${\mathbb{E}}_{{\mathcal{D}}}\Big[\sup_{f\in{\mathscr{F}}}R(\widehat{f})-2\widehat{R}_{n}(\widehat{f})\Big]\leq{\mathbb{E}}_{{\mathcal{D}}}{\mathbb{E}}_{\xi}\Big[\sup_{h\in{\mathscr{H}}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h^{2}(X^{(i)})\Big].$
We next estimate the expectation in the right-hand side of (I.1). Let $\delta\in(0,4B^{2})$ and ${\mathscr{H}}_{\delta}$ be a $L^{\infty}({\mathcal{D}})$ $\delta$-cover of $\mathscr{H}$} satisfying $|{\mathscr{H}}_{\delta}|=N(\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))$. Then for each $h\in{\mathscr{H}}$, there exists $h_{\delta}\in{\mathscr{H}}_{\delta}$ such that
@@
$\frac{1}{n}\sum_{i=1}^{n}h^{2}(X^{(i)})+\frac{1}{n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\leq 8B^{2}\delta.$
@@
(I.2) $\leq{\mathbb{E}}_{\xi}\Big[\max_{h_{\delta}\in{\mathscr{H}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\Big]+5\delta.$
@@
$h_{\delta}(X^{(i)})\leq\xi^{(i)}h_{\delta}(X^{(i)})\leq h_{\delta}(X^{(i)}),\quad 1\leq i\leq n.$
@@
(I.3) $\leq\exp\Big(-\frac{(\frac{nt}{3}+\frac{1}{12B^{2}}\sum_{i=1}^{n}h_{\delta}^{2}(t^{(i)},X_{t}^{(i)}))^{2}}{2\sum_{i=1}^{n}h_{\delta}^{2}(t^{(i)},X_{t}^{(i)})}\Big)\leq\exp\Big(-\frac{nt}{18B^{2}}\Big),$
@@
where the inequality is owing to (I.3). Letting $A=18B^{2}\log|{\mathscr{H}}_{\delta}|n^{-1}$ gives that
(I.4) ${\mathbb{E}}_{\xi}\Big[\max_{h_{\delta}\in{\mathscr{H}}_{\delta}}\frac{3}{n}\sum_{i=1}^{n}\xi^{(i)}h_{\delta}(X^{(i)})-\frac{1}{4B^{2}n}\sum_{i=1}^{n}h_{\delta}^{2}(X^{(i)})\Big]\leq 18B^{2}\frac{\log|{\mathscr{H}}_{\delta}|+1}{n}.$
@@
(I.5) $\log N(\delta,{\mathscr{H}},L^{\infty}({\mathcal{D}}))\leq\log N\Big(\frac{\delta}{4B},{\mathscr{F}},L^{\infty}({\mathcal{D}})\Big)\leq\operatorname{VCdim}({\mathscr{F}})\log\Big(\frac{4eB^{2}n}{\delta}\Big),$
where the first and last inequalities follows from Lemmas B.10 and B.11, respectively. Combining (I.1), (I.2), (I.4) and (I.5), we have
[…]
@@
Lemma I.2.
Suppose that $|f^{*}(x)|\leq B$ for each $x\in{\mathcal{X}}$. Let $\mathscr{F}$} be a set of functions mapping from $\mathcal{X}$} to [B,B][-B,B]. Let $\{\varepsilon^{(i)}\}_{i=1}^{n}$ be a set of independent $\sigma^{2}$-sub-Gaussian random variables. Then it follows that for each $n\geq\operatorname{VCdim}({\mathscr{F}})$,
@@
This proof uses a technique similar to the proof of [Schmidt-Hieber, 2020, Lemma 4].
Proof of Lemma I.2.
@@
(I.6) ${\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big[\frac{1}{n}\sum_{i=1}^{n}\varepsilon^{(i)}(\widehat{f}(X^{(i)})-\widehat{f}_{\delta}(X^{(i)}))\Big]\leq\delta{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big[\frac{1}{n}\sum_{i=1}^{n}|\varepsilon^{(i)}|\Big]\leq\delta\sigma,$
@@
(I.7) $\widehat{R}_{n}^{1/2}(\widehat{f}_{\delta})-\widehat{R}_{n}^{1/2}(\widehat{f})\leq\Big(\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{\delta}(X^{(i)})-\widehat{f}(X^{(i)}))^{2}\Big)^{1/2}\leq\delta.$
@@
(I.8) $\leq\frac{1}{4}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\Big[\widehat{R}_{n}(\widehat{f})\Big]+\frac{2}{n}{\mathbb{E}}_{({\mathcal{D}},\varepsilon)}\big[\psi^{2}(\widehat{f}_{\delta})\big]+\frac{1}{4}\delta^{2}+\delta\sigma,$
@@
Here the first inequality follows from (I.6), the second inequality holds from (I.7), the third inequality is due to Cauchy-Schwarz inequality, and the last inequality is owing to the weighted AM-GM inequality $ab\leq a/4+b$ for each $a,b\in{\mathbb{R}}$.
@@
(I.9) ${\mathbb{E}}_{\varepsilon}\big[\psi^{2}(\widehat{f}_{\delta})\big]\leq{\mathbb{E}}_{\varepsilon}\Big[\max_{f_{\delta}\in{\mathscr{F}}_{\delta}}\psi^{2}(f_{\delta})\Big]\leq 4\sigma^{2}(\log|{\mathscr{F}}_{\delta}|+1).$
We now estimate the covering number $|{\mathscr{F}}_{\delta}|=N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))$. It follows from Lemma B.11 that for $n\geq\operatorname{VCdim}({\mathscr{F}})$,
(I.10) $\log N(\delta,{\mathscr{F}},L^{\infty}({\mathcal{D}}))\leq\operatorname{VCdim}({\mathscr{F}})\log\Big(\frac{eBn}{\delta}\Big).$
Combining (I.8), (I.9) and (I.10), and setting $\delta=B/n$ complete the proof. ∎
Appendix J Approximation by Deep Neural Networks with Lipschitz Constraint
The approximation error analysis for deep neural networks has been investigated by Yarotsky [2017, 2018], Yarotsky and Zhevnerchuk [2020], Shen et al. [2019], Shen [2020], Lu et al. [2021], Petersen and Voigtlaender [2018], Jiao et al. [2023a], Duan et al. [2022]. However, limited work has been done for deep neural networks with Lipschitz constraint Huang et al. [2022], Chen et al. [2022], Jiao et al. [2023b], Ding et al. [2024].
This proof is based on the proof of [Yarotsky, 2017, Theorem 1]. The ReLU activation function is defined as $\operatorname{ReLU}(x)=\max\{0,x\}$, and the ReQU activation function is defined as the squared ReLU function $\operatorname{ReQU}(x)=(\max\{0,x\})^{2}$.
Lemma J.1.
@@
Proof of Lemma J.2.
According to [Li et al., 2019, Lemma 2.1], the following identities hold:
@@
Proof of Lemma J.3.
Define the augmented input vector $(x_{1},\ldots,x_{p},x_{p+1},\ldots,x_{n})$ where $n=2^{\lceil\log_{2}p\rceil}$ and $x_{i}=1$ for $p+1\leq i\leq n$. Observe that $\prod_{i=1}^{p}x_{i}=\prod_{i=1}^{n}x_{i}$. According to Lemma J.2, the mapping $(x_{1},\ldots,x_{n})\rightarrow(x_{1}x_{2},\ldots,x_{n-1}x_{n})\in{\mathbb{R}}^{n/2}$ can be implemented by a ReQU neural network with $1$ hidden layer and $6n$ non-zero parameters. By a same argument, we can construct a ReQU neural network with $1$ hidden layer and $3n$ non-zero parameters, which maps $(x_{1}x_{2},\ldots,x_{n-1}x_{n})$ to $(x_{1}x_{2}x_{3}x_{4},\ldots,x_{n-3}x_{n-2}x_{n-1}x_{n})\in{\mathbb{R}}^{n/4}$. By induction on $n$, the number of layers is given as $\lceil\log_{2}p\rceil$, and the total number of non-zero parameters is given by
@@
(J.1) $\psi(z)=\left\{\begin{aligned} &1,&&|z|<1,\\ &2-|z|,&&1\leq|z|\leq 2,\\ &0,&&2<|z|,\end{aligned}\right$.
@@
Noticing that $\widetilde{\psi}(z)=\min\{\operatorname{ReLU}(z+2),\operatorname{ReLU}(-z+2)\}$, by applying Lemma J.1, we find that $\widetilde{\psi}$ can be implemented by a ReLU neural network with $2$ hidden layers and $11$ non-zero parameters. Further, according to the equality $\psi(z)=\min\{1,\widetilde{\psi}(z)\}$, the univariate trapezoid function $\psi$ can be implemented by a ReLU neural network with $3$ hidden layers and $14$ non-zero parameters. This completes the proof. ∎
Lemma J.5 (Approximation error).
@@
Proof of Lemma J.5.
The proof is divided into three steps. In the first step, we approximate the target function based on a partition of unity and the degree-0 Taylor expansion. Then we implement this piece-wise linear function using deep neural network exactly in the second step. Finally, in the last step, we estimate the Lipschitz constant of the deep neural network.
@@
Consider a partition of unity formed by a grid of $\prod_{k=1}^{p}(N_{k}+1)$ functions $\phi_{m}$ on the domain $0,1]^{p}$:
(J.2) $\sum_{m}\phi_{m}(x)\equiv 1,\quad x\in[0,1]^{p},$
@@
(J.3) $\phi_{m}(x)=\prod_{k=1}^{p}\psi\Big(3N_{k}\Big(x_{k}-\frac{m_{k}}{N_{k}}\Big)\Big).$
Here $\psi$ is the univariate trapezoid function defined as (J.1). It is noticeable that for each $m$,
(J.4) $\sup_{z\in[0,1]}|\psi(z)|=1,\quad\sup_{x\in[0,1]^{p}}|\phi_{m}(x)|=1,$
@@
(J.5) $\operatorname{supp}(\phi_{m})\subseteq\Big\{x\in[0,1]^{p}:\Big|x_{k}-\frac{m_{k}}{N_{k}}\Big|\leq\frac{2}{3N_{k}},\penalty 10000\ 1\leq k\leq p\Big\}.$
@@
(J.6) $u(x)=\sum_{m}\phi_{m}(x)u^{*}\Big(\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big).$
@@
$\leq\sum_{m}\Big|u^{*}(x)-u^{*}\Big(\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big)\Big|\mathbbm{1}\Big\{m:\Big|x_{k}-\frac{m_{k}}{N_{k}}\Big|\leq\frac{2}{3N_{k}},\penalty 10000\ k\in[p]\Big\}$
$\leq 2^{p}\max_{m}\Big|u^{*}(x)-u^{*}\Big(\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big)\Big|\mathbbm{1}\Big\{m:\Big|x_{k}-\frac{m_{k}}{N_{k}}\Big|\leq\frac{2}{3N_{k}},\penalty 10000\ k\in[p]\Big\}$
$\leq 2^{p}\max_{m}\sum_{k=1}^{p}\mathop{\operatorname{ess}\,\sup}_{x\in[0,1]^{p}}|\partial_{k}u^{*}(x)|\Big|x_{k}-\frac{m_{k}}{N_{k}}\Big|\mathbbm{1}\Big\{m:\Big|x_{k}-\frac{m_{k}}{N_{k}}\Big|\leq\frac{2}{3N_{k}},\penalty 10000\ k\in[p]\Big\}$
@@
where the first inequality follows from the triangular inequality, the second inequality is due to (J.4) and (J.5), the third inequality used the observation that any $x\in[0,1]^{p}$ belongs to the support of at most $2^{d}$ functions $\phi_{m}$, the forth inequality used Taylor’s theorem of degree-0, and the last inequality holds from Hölder’s inequality. Consequently, we obtain the following inequality
@@
In this step, we implement the piececwise linear approximation (J.6) by a deep neural network. Using Lemmas J.3 and J.4, for each $m$, the function $\phi_{m}$ defined as (J.3) can be implemented by a deep neural network with $\lceil\log_{2}p\rceil+3$ layers and $16p+6(p+1)=22p+6$ non-zero parameters. Since $u$ defined in (J.6) is a linear combination of $\prod_{k=1}^{p}(N_{k}+1)$ functions $\phi_{m}$, it can be implemented by a deep neural network with $\lceil\log_{2}p\rceil+3$ layers and $(22p+6)\prod_{k=1}^{p}(N_{k}+1)$ non-zero parameters.
@@
According to Step 2, the piecewise linear approximation $u$ can be implemented by a deep neural network with no error. Therefore, it suffices to compute the Lipschitz constant of $u$ in (J.6). Taking derivative on both sides of (J.6) with respect to $x_{k}$ yields
@@
(J.7) $=\sum_{m}u^{*}\Big(\frac{m_{1}}{N_{1}},\ldots,\frac{m_{p}}{N_{p}}\Big)A_{k}(m)\partial_{k}\psi\Big(3N_{k}\Big(x_{k}-\frac{m_{k}}{N_{k}}\Big)\Big)\mathbbm{1}\Big\{m:\Big|x_{k}-\frac{m_{k}}{N_{k}}\Big|\leq\frac{2}{3N_{k}}\Big\},$